OpenStudy (anonymous):

Which set of measurements would NOT form a triangle? 13, 4, 8 14, 18, 13 13, 8, 14 18, 14, 8

OpenStudy (ivettef365):

a+b >c a+c>b b+c>a

OpenStudy (anonymous):

?

OpenStudy (ivettef365):

ok, to see which one is a triangle add any 2 sides should be more than the third side

OpenStudy (ivettef365):

example, lets try 13, 8, 14 13 + 8 = 21 > 14 is ok 13 + 14 = 27 > 8 is ok 8 + 14 = 22 > 13 is ok

OpenStudy (ivettef365):

now lets try 13, 4, 8 13 + 4 = 17 > 8 is ok 13 + 8 = 21 > 13 is ok 4 + 8 = 12 > 13 NOT TRUE, this can not be a triangle

OpenStudy (anonymous):

Ohh i understand now. Thanks

OpenStudy (ivettef365):

yw :)